OpenStudy (itsbribro):

What are the coordinates of the circumcenter of a triangle with vertices A(−3,−4), B(1,−4), and C(−3,0)?

OpenStudy (itsbribro):

idk how to do this

OpenStudy (tkhunny):

You will need Perpendicular bisectors on two sides.

OpenStudy (itsbribro):

is it one of the points given?

OpenStudy (itsbribro):

@eliesaab

OpenStudy (tkhunny):

No, those are the vertices of the triangle. Use to pair to find the perpendicular bisector of the included side. Example: A (-3,-4) vs B (1,-4) Slope of AB: (-4+4)/(1+3) = 0/4 = 0 -- Oh, this is a horizontal line. Midpoint of AB: ((-3+1)/2,(-4-4)/2) = (-2/2,-8/2) = (-1,-4) Slope of Perpendicular Bisector is "No Slope". It's a vertical line Circumcenter is SOMEWHERE on the line x = -1. You do the same thing with A vs C. Note: Plot the points. See for yourself that using BC is going to be harder than using AC.

OpenStudy (itsbribro):

thank you

OpenStudy (tkhunny):

It's a little harder if you don't get horizontal or vertical lines. Same idea, though.

OpenStudy (itsbribro):

yeah